# SP2T: Sparse Proxy Attention for Dual-stream Point Transformer

### v1

@@ Abstract
In 3D understanding, point transformers have yielded significant advances in broadening the receptive field. However, further enhancement of the receptive field is hindered by the constraints of grouping attention. The proxy-based model, as a hot topic in image and language feature extraction, uses global or local proxies to expand the model’s receptive field. But global proxy-based methods fail to precisely determine proxy positions and are not suited for tasks like segmentation and detection in the point cloud, and exist local proxy-based methods for image face difficulties in global-local balance, proxy sampling in various point clouds, and parallel cross-attention computation for sparse association. In this paper, we present SP2T, a local proxy-based dual stream point transformer, which promotes global receptive field while maintaining a balance between local and global information. To tackle robust 3D proxy sampling, we propose a spatial-wise proxy sampling with vertex-based point proxy associations, ensuring robust point-cloud sampling in many scales of point cloud. To resolve economical association computation, we introduce sparse proxy attention combined with table-based relative bias, which enables low-cost and precise interactions between proxy and point features. Comprehensive experiments across multiple datasets reveal that our model achieves SOTA performance in downstream tasks. The code has been released in this url.
@@ 1 Introduction
The point transformer [50, 43, 28, 44] has recently emerged as a key research area in 3D perception [42] and other field [48], mainly using group attention [24, 25] to extract features. Although various point-transformer methods [43, 44] have recently advanced in expanding receptive fields, they remain constrained by the immense scale of point-cloud data. For example, the receptive field of the Point Transformer has expanded from 16 points in PTv2 [43] to 1024 points in PTv3 [44]. However, the total number of point clouds can range from 10,000 to 100,000 [10, 1], as shown in Fig. 1. Due to the constraints of group attention, further enlarging the receptive field in point transformer could be counterproductive. Thus, the receptive field must be broadened through alternative approaches.
For the extraction of language and image features, the proxy-based methods [41, 18, 17, 16, 11, 14] have emerged as a focus area, expanding the receptive field of neural networks while demanding less computation. These methods are generally divided into global proxy-based methods [41, 18, 17, 11] and local proxy-based methods [22, 16, 14]. Global proxy-based methods utilize proxies to exchange global information, thereby expanding the model’s receptive field. However, the methods do not accurately account for the exact positioning of the proxies and are less suited to tasks focused on local details such as segmentation and detection. Conversely, local proxy-based methods utilize the proxy sampling method to determine the proxy’s position and apply the association to depict the relationship between data and proxy, which expand the receptive field while ensure feature localization. However, the sampling and associations method for images [16, 14] proves to be challenging to implement for point clouds that exhibit diverse scales and size proportions. Moreover, when dealing with sparse point clouds, effectively parallelizing the association process based on cross-attention between proxies and points poses challenges in terms of time complexity and memory usage.
Figure 1: Receptive field comparison between PTv3 (Light Red) and ours (Overall Color).
Furthermore, networks that employ proxy methods must strike a balance between global and local information. In tasks like segmentation or detection, both global and local information is of equal significance. Global features with contextual information, on the one hand, improve classification results. Conversely, the feature of the local point cloud are crucial for refining segmentation and detection bounding boxes. Drawing inspiration from MobileFormer [5], we propose that a dual-stream Transformer architecture aids in harmonizing local and global information extraction.
To address the aforementioned issues, a local proxy-based dual-stream transformer, named SP2T is introduced for point cloud understanding. Our approach involves a dual-stream Point Transformer to address the challenge of integrating local details and global information in point clouds. In SP2T, both the proxy and the point cloud are fed into the network simultaneously, allowing the balance extraction of global and local information.
Secondly, concerning proxy sampling in point clouds, we introduce a spatial-wise method employing a binary selection approach to determine the optimal proxy distribution. Subsequently, we propose a vertex-based point-proxy association method that speeds up the establishment of point-proxy associations by examining coordinate relationships through spatial-based proxy sampling.
Thirdly, to address the difficulty of efficiently computing cross-attention between the sparsely connected proxy and point, we present the sparse proxy attention (SPA). Specifically, SPA has changed from using the matrix-based attention method to using association lists, and it uses Map-Reduce [8] for parallel speed-up. Furthermore, a Table-Based Relative Bias is introduced to improve spatial awareness of SPA.
In general, the contributions of this paper include the following parts: (1) We introduce SP2T, a local proxy-based dual-stream point transformer, which effectively facilitates global receptive field while preserving a balance between global and local. (2) For the proxy sampling and association method in the sparse point cloud, a spatial-wise proxy sampling method with a vertex-based association is introduced, which ensures efficient and reliable sampling of the proxies and calculation of the association in the point cloud. (3) To address the issue of efficiently transferring information between points and proxies, Sparse Proxy Attention is introduced along with table-based relative bias. (4) Extensive experiments conducted on the many datasets [10, 1, 3] demonstrate that our model achieves state-of-the-art (SOTA) results in many downstream tasks.
@@ 2 Related Works
3D Point Cloud Understanding. Currently, methods oriented towards 3D Point Cloud Understanding can be divided into three main categories: 2D projection-based [36, 23, 4, 21], voxel-based [26, 34, 51, 6, 12, 39, 47, 15], and point cloud-based methods [29, 30, 49, 31, 50, 43, 44]. Initially, the 2D projection method utilizes the project to derive features from the point cloud using a 2D-backbone, and the voxel-based approach converts the point cloud into a structured voxel grid, which speeds up convolution or attention computations. Moreover, the point-cloud-based method processes the point-cloud directly using pooling-based or transformer architectures. SP2T is classified as a point-cloud-based network, which employs proxies to significantly expand the receptive field of the point-cloud network, thereby enhancing the model’s proficiency in understanding 3D point clouds.
Proxy-based Methods. Recently, proxy-based methods for feature extraction in both imagery and linguistics have gained considerable attention. These methods are categorized mainly into global proxy-based approaches [41, 18, 17, 11] and local proxy-based approaches [22, 16, 14]. Initially, global proxy-based techniques employ proxies to gather general information, enhancing the model’s receptive field. However, global proxies lack explicit modeling of proxy location and are not friendly to local-relevant tasks such as segmentation and detection. Secondly, local proxy-based methods focus solely on aggregating features around the proxy and incorporating information exchange between agents, thereby achieving the global receptive field. For the local proxy-based method, the sampling method and association strategy of the proxy are very important, which determine the spatial location of the proxy and the extraction range of the image. However, the sampling and association of the local proxy for point cloud data still require further investigation. SP2T employs a local proxy-based framework and examines the sampling and association method for point clouds, utilizing sparse proxy attention to expedite the computation of proxy-point interactions.
@@ 3.1 Overall Architecture
SP2T. As illustrated in Fig. 2, SP2T is a dual-stream encoder-decoder network specifically designed for point-cloud data. Formally, the input points $P_{point}\in\mathbb{R}^{N\times 3}$ are processed through the point embedding module and the proxy initialization module to derive the feature of the point $F_{point}\in\mathbb{R}^{N\times C}$ and the proxy $F_{proxy}\in\mathbb{R}^{M\times C}$ and the position of the proxy $P_{point}\in\mathbb{R}^{M\times 3}$, where $N$ and $M$ is the number of points and proxy, and $C$ is the embed channel.
Afterward, the positions and features of both the points and proxies are fed into a multi-stage encoder-decoder network. Each SP2TLayer is made up of several SP2TBlock for the exaction of features and ends with a downsampling or upsampling module for the point cloud and proxy to facilitate the encoding and decoding processes. Finally, the features produced by the last layer of the decoder are utilized for downstream tasks. For detailed network parameters, we report them in the support material.
SP2TBlock. The SP2TBlock is the foundational module in SP2T, enabling distinct local feature aggregation for points and global feature aggregation for proxies. SP2TBlock contains four distinct modules for feature extraction: local fusion, global fusion, point-proxy interaction, and proxy-point interaction module.
Specifically, the local fusion module is crafted to seamlessly integrate the local features of point-cloud data, while the global fusion module exceeds local receptive field limitations for a comprehensive global context understanding. Subsequently, the point-proxy and proxy-point interaction modules combine information from points and proxies using the sparse proxy attention (SPA). These modules primarily enhance the exchange of information between the point cloud and proxy, enriching point cloud features with detailed semantic context while maintaining the integrity of the proxies’ local feature representation.
@@ 3.2 Proxy Sampling and Association
Problem of FPS-based Proxy Sampling. FPS-based proxy sampling utilizes the Furthest Point Sampling algorithm [30] to select a constant number of reference points from the point cloud, as widely applied in PointNet[29] and other methods. But in our method, the FPS-based approach results in a decrease in precision. Using FPS-based sampling generates a scene-relevant proxy. The model might overfit to such proxy distribution during training, resulting in a segmentation output derived directly from the scene-relevant proxy rather than from the point feature.
Problem of Grid-based Proxy Sampling Regarding grid-based proxy sampling methods, there are two primary approaches to sampling: one maintains a constant number of proxies, while the other regulates a specific proxy spacing. However, each of these methods has certain challenges. Firstly, the sampling method that keeps the number of proxies constant uses a predetermined quantity $X_{voxel}\times Y_{voxel}\times Z_{voxel}$ within a uniform range. However, as illustrated in Fig. 3 (a), since the aspect ratio of the generated proxies mirrors that of the axis-aligned bounding box (AABB) of the point cloud, the proxies will appear stretched if the AABB has an extreme aspect ratio. This uneven distribution along different directions can adversely affect the uniformity of associations and feature fusion.
Secondly, the spacing-fixed sampling method separates the aspect ratio of the proxy’s spacing from the point cloud by fixing the size of the proxy’s spacing, enabling the uniformity of the proxies’ distribution in space. However, as illustrated in Fig. 3 (b), since the AABB’s range of the point cloud can fluctuate, the resultant number of proxies generated also varies, introducing significant uncertainty regarding the FLOPs of the model.
@@
Spatial-wise Proxy Sampling. To address the aforementioned issues, we propose a spatial-wise proxy sampling method. Specifically, we introduced a binary search method for searching the proxy spacing. It is noted that, under an equal proxy spacing along different axis, the number of generated proxies monotonically decreases as the proxy spacing increases. Therefore, given a desired range of proxy numbers, the binary search method can be used to find an appropriate proxy spacing that brings the number of generated proxies as close as possible to the required range.
In practice, for a fixed number of proxies, our sampling method efficiently discerns the ideal proxy spacing by considering the AABB sizes of various point clouds. The method is designed to maintain the count of proxies within the bounds of $N_{max}$ and $N_{min}$, using a bisection approach to determine the optimal proxy spacing $L_{p}$. If the number of proxies exceeds $N_{max}$, the proxy spacing $L_{p}$ is reduced and the total number of proxies is recalculated. In contrast, $L_{p}$ increases until the number of proxies reaches the desired range from $N_{max}$ to $N_{min}$. As demonstrated in Fig. 3 (c), this method autonomously selects the best proxy spacing based on the predetermined proxy count for different point clouds depending on their AABB size. Compared to the other sampling methods in Fig. 3 (a) and (b), the spatial-wise proxy sampling method guarantees an appropriate aspect ratio for the proxy spacing and specifies a distinct number of proxies.
Vertex-based Association. A commonly adopted method for association uses the K-Nearest Neighbor (KNN) [37] method, which connects each data point with its $k$ nearest proxies. This approach provides considerable flexibility in defining the degree of correlation between the proxies and the point cloud, thereby permitting modifications in the number of associations.
Despite this, the KNN method requires evaluating the distances between all point clouds and proxies, which adds computational complexity. In KNN using the L-Infinity Norm with $k=8$, every point is linked to eight proxies positioned at the corners of a cubic proxy within the AABB coordinate sphere. As the proxy is formed through grid sampling, its structural layout can be exploited to create a more efficient association method, thereby reducing the need for intricate distance computations.
As illustrated in Fig. 4, the initial transformation involves converting the coordinates of the points to their corresponding locations within the proxy grid. Then, by computing the floor and ceiling for each coordinate, the positions of the eight nearest proxies are determined. Finally, the compilation of association tuples is achieved by aggregating all associations into one comprehensive list.

### v2

@@ Abstract
Point transformers have demonstrated remarkable progress in 3D understanding through expanded receptive fields (RF), but further expanding the RF leads to dilution in group attention and decreases detailed feature extraction capability. Proxy, which serves as abstract representations for simplifying feature maps, enables global RF. However, existing proxy-based approaches face critical limitations: Global proxies incur quadratic complexity for large-scale point clouds and suffer positional ambiguity, while local proxy alternatives struggle with 1) Unreliable sampling from the geometrically diverse point cloud, 2) Inefficient proxy interaction computation, and 3) Imbalanced local-global information fusion; To address these challenges, we propose Sparse Proxy Point Transformer (SP2T) – a local proxy-based dual-stream point transformer with three key innovations: First, for reliable sampling, spatial-wise proxy sampling with vertex-based associations enables robust sampling on geometrically diverse point clouds. Second, for efficient proxy interaction, sparse proxy attention with a table-based relative bias effectively achieves the interaction with efficient map-reduce computation. Third, for local-global information fusion, our dual-stream architecture maintains local-global balance through parallel branches. Comprehensive experiments reveal that SP2T sets state-of-the-art results with acceptable latency on indoor and outdoor 3D comprehension benchmarks, demonstrating marked improvement (+3.8% mIoU vs. SPoTr@S3DIS, +22.9% mIoU vs. PointASNL@Sem.KITTI) compared to other proxy-based point cloud methods.
@@ 1 Introduction
The point transformer [68, 55, 37, 56] has recently emerged as a key research area in 3D perception [54, 17, 29] and other fields [63], mainly using group attention [31, 32] to extract features. Although various point transformer methods [55, 56] have recently advanced in expanding receptive fields (RF), they remain constrained by the large scale of points. However, expanding RF within PTv3 can lead to attention dilution [39], where a larger group size diverts the model’s focus, impairing its ability to capture precise features (77.3 mIOU with 1024 points vs. 77.1 mIOU with 4096 points@ScanNetv2 [11]). Furthermore, this expansion is also computationally demanding. Consequently, the broadening of the RF must be achieved differently.
Figure 1: The comparison with other Proxy-based Methods. The improvement of SP2T compared with other proxy-based method are threefold: (1) Advances Sampling, (2) Efficiency Interaction and (3) Enhanced Dual-stream Framework.
In recent years, proxies, which serve as key elements on the feature map or represent an abstraction of features for simplification, have attracted significant interest from the research community. Proxy-based methods are typically classified into global [53, 21, 20, 12, 21] and local approaches [27, 19, 15, 57, 38], both with a global RF. However, global proxy methods have quadratic complexity, and the positional uncertainty of proxies adversely affects segmentation and detection tasks such as object boundary localization, making them unsuitable for point feature extraction.
In contrast, local proxy-based methods [27, 19, 15, 57, 38] employ a proxy sampling technique to establish the location of the proxy. They utilized the association to illustrate the link between data and proxy, with low-cost complexity and precise positioning ideal for extracting the feature of points. However, existing local proxy-based methods face challenges such as 1) inconsistent sampling from geometrically diverse points, 2) inefficient computation of proxy interactions, and 3) imbalanced integration of local and global information, as shown in Fig. 1. Initially, the sampling and association methods outlined are challenging to implement in geometrically diverse points, leading to an unstable spatial range of proxy. Secondly, handling sparse points introduces difficulties in optimizing the association process with cross-attention between proxies and points, particularly in terms of complexity and memory requirements with a considerable number of points. Third, existing proxy methods for points such as SPoTr [38] and Fast PT [37] lose the balance between global and local information. In tasks like segmentation, global features are critical for class-level understanding, while local features define object boundaries; thus, striking a balance is crucial.
To address these challenges, we propose the Sparse Proxy Point Transformer (SP2 T), a dual stream transformer with local proxies to understand the global receptive field. Inspired by MobileFormer [6], PointASNL [57], and others [34, 66], SP2T employs a dual stream structure: a proxy stream for compressed global representations and a point stream for detailed geometric features.
Secondly, for robust proxy sampling, we propose a spatial-wise binary selection method for an optimal proxy distribution. Compared to FPS, it reduces complexity, ensures uniformity, and mitigates information leakage from aggregation at the farthest point. We also introduce vertex-based association, aligning points to grid-vertex proxies equivalent to L-infinity KNN, enabling rapid performance.
Third, we propose Sparse Proxy Attention (SPA) for efficient cross-attention between sparsely connected proxies and points. SPA replaces matrix-based attention with linear complexity and employs Map-Reduce for parallel speedup. Additionally, Table-Based Relative Bias (TRB) is introduced to overcome traditional Relative Bias’ limitations in sparse data, enhancing SPA’s spatial perception.
In general, our improvement compared with other proxy-based method are following: (1) Advances Sampling: Replaces FPS and learning-based method with a spatial-wise method using vertex-based association, achieving explainability and scenario adaptability through various 3D scenes. (2) Efficiency Interaction: SPA reduces the computation of $O(nmd)$ of dense attention to $O(nkd+m^{2}d)(k=8,m<<n)$, facilitating the implementation of proxy-based methods on large 3D scenes. (3) Enhanced Dual-stream Framework: Optimized with attention-based local fusion to improve local extraction with global-local equilibrium. (4) SOTA Performance: Extensive experiments indicate that SP2T achieves state-of-the-art performance with acceptable latency (20 FPS in nuScenes [4]) on indoor and outdoor 3D understanding benchmarks [11, 43, 1, 4, 2, 48], showing significant improvements (+3.8% mIoU compared to SPoTr@S3DIS [1], +22.9% mIoU compared to PointASNL@Sem.KITTI [2]) relative to other proxy-based point cloud methods.
@@ 2 Related Works
3D Point Cloud Understanding. Currently, methods oriented towards 3D Point Cloud Understanding can be divided into three main categories: 2D projection-based [47, 28, 5, 26], voxel-based [35, 45, 69, 7, 13, 51, 62, 18], and point cloud-based methods [40, 41, 67, 42, 68, 55, 56, 30] for down-stream task [64, 33, 65]. SP2T is classified as a point-cloud-based network, which uses proxies to expand the receptive field of the point-cloud network significantly. Compared to PTv3 [56], the primary focus of PTv3 is serialization attention, which is a unique style of group attention [30]. However, further expansion of the window size of serialization attention diverts the model’s focus, impairing its ability to capture precise features. In contrast, SPA uses proxies to maintain feature perception with a global receptive field. Additionally, our hybrid structure takes advantage of PTv3’s serialization attention for capturing local patterns and SPA for facilitating global interactions.
Proxy-based Methods. Recently, proxy-based methods for feature extraction have gained considerable attention. These methods are categorized mainly into global proxy-based approaches [53, 21, 20, 12, 37, 21] and local proxy-based approaches [27, 19, 15, 57, 38]. The proxy method faces challenges such as complexity, sampling, and association methods. Consider Register ViT [12] (global) and STViT [19] (local) as case studies. The ViT register faces a complexity $O(n^{2})$, leading to inefficiencies with large data sets, while the STViT scratch window sampling, intended for 2D data, is not well equipped to handle the sparse nature of point clouds. Furthermore, PointASNL [57], as a proxy-based point network, is severely limited to a crude point-proxy interaction design. In contrast, SP2T employs a local proxy-based framework and examines the spatial-wise sampling and vertex-based association method for point clouds, using sparse proxy attention to accelerate the computation of proxy-point interactions. Sparse Proxy Attention efficiently establishes sparse associations between points and proxies, making it better suited for 3D environments. In addition, some proxy-based models for point clouds, such as Fast PT [37] and SPoTr [38], overlook the equilibrium between local and global information, resulting in poor performance in indoor datasets. SP2T employs a network architecture that harmonizes global and local information, enabling proxy-based models to achieve competitive precision in outdoor and indoor data sets for the first time.
@@ 3.1 Overall Architecture
SP2T. As illustrated in Fig. 2, SP2T is a dual-stream network specifically designed for points. Formally, the input points $P_{point}\in\mathbb{R}^{N\times 3}$ are processed through the point embedding module and the proxy initialization module to derive the feature of the point $F_{point}^{0}\in\mathbb{R}^{N\times C}$ and the proxy $F_{proxy}^{0}\in\mathbb{R}^{M\times C}$ and the position of the proxy $P_{proxy}\in\mathbb{R}^{M\times 3}$, where $N$ and $M$ is the number of points and proxy, and $C$ is the channel. The point embedding comprises several MLPs, while the proxy initialization component includes sampling, association, and several MLPs to synthesize $F_{proxy}^{0}$ from points associated with proxies. For the proxy feature $F_{Proxy}$, the model applies sinusoidal position embedding to define its feature as $F_{Proxy}=\text{Sinusoidal}(P_{proxy})$ after the proxy identifies its location $P_{proxy}\in\mathbb{R}^{3}$ using proxy sampling.
Afterward, the positions and features of both the points and proxies are fed into a multi-stage encoder-decoder network. Each SP2TLayer is made up of several SP2TBlock for the exaction of features and ends with an up/downsampling module for points $P_{point}^{i},F_{point}^{i}$ and a proxy $P_{proxy}^{i},F_{proxy}^{i}$ to facilitate the encoding and decoding processes, where $i$ is the idx of Each SP2TLayer. Finally, the features and location of the points $P_{point}^{k},F_{point}^{k}$ produced by the last layer $k$ of the decoder are utilized for downstream tasks. For network details, we report them in supplementary materials.
SP2TBlock. The SP2TBlock is the foundational module in SP2T, enabling distinct local feature aggregation for points and global feature aggregation for proxies. SP2TBlock contains four distinct modules for feature extraction: local fusion, global fusion, point-proxy interaction, and proxy-point interaction module. Specifically, the local fusion module of the layer $i$ is designed to seamlessly integrate the local feature of the points $F_{proxy}^{i}$. In contrast, the global fusion module exceeds the limitations of the local receptive field for a comprehensive understanding of the global context. Subsequently, the point-proxy and proxy-point interaction modules combine the features of the points with the proxies $F_{proxy}^{i}$ using sparse proxy attention (SPA). These modules primarily enhance the exchange of information between the points and the proxy, enriching the features of the points with a detailed semantic context while maintaining the integrity of the local representation of the features of the proxies.
@@ 3.2 Proxy Sampling and Association
Problem of FPS-based Proxy Sampling. FPS-based proxy sampling utilizes the Furthest Point Sampling algorithm [41] to select a constant number of reference points from the points, as widely applied in PointNet [40] and other methods. However, in our method, the FPS-based approach decreases precision. FPS-based sampling generates a scene-relevant proxy. The model might overfit such a proxy distribution during training, resulting in a segmentation output derived directly from the scene-relevant proxy rather than the point feature.
Problem of Grid-based Proxy Sampling. Regarding grid-based proxy sampling methods, there are two primary approaches to sampling: one maintains a constant number of proxies, while the other regulates a specific proxy spacing. However, each of these methods has particular challenges. Firstly, the sampling method that keeps the number of proxies constant uses a predetermined quantity $X_{voxel}\times Y_{voxel}\times Z_{voxel}$ within a uniform range. However, as illustrated in Fig. 3 (a), since the aspect ratio of the generated proxies mirrors that of the axis-aligned bounding box (AABB) of the points, the proxies will appear stretched if the AABB has an extreme aspect ratio. This uneven distribution along different directions can adversely affect the uniformity of associations and feature fusion. Secondly, the spacing-fixed sampling method separates the aspect ratio of the proxy’s spacing from the points by fixing the size of the proxy’s spacing, enabling the uniformity of the proxies’ distribution in space. However, as illustrated in Fig. 3 (b), since the AABB’s range of the points can fluctuate, the resultant number of proxies generated also varies, introducing significant uncertainty regarding the FLOPs of the model.
@@
Spatial-wise Proxy Sampling. We propose a spatial-wise proxy sampling method to address the issues mentioned above. Specifically, we introduced a binary search method for finding the proxy spacing. It is noted that, under equal proxy spacing along different axes, the number of generated proxies monotonically decreases as the proxy spacing increases. Therefore, given a desired range of proxy numbers, the binary search method can be used to find an appropriate proxy spacing that brings the number of generated proxies as close as possible to the required range.
In practice, our sampling method efficiently discerns the ideal proxy spacing for a fixed number of proxies by considering the AABB sizes of various points. The technique is designed to maintain the proxy count within the bounds of $N_{max}$ and $N_{min}$, using a bisection approach to determine the optimal proxy spacing $L_{p}$. If the number of proxies exceeds $N_{max}$, the proxy spacing $L_{p}$ is reduced, and the total number of proxies is recalculated. In contrast, $L_{p}$ increases until the number of proxies reaches the desired range from $N_{max}$ to $N_{min}$. As shown in Fig. 3 (c), this method autonomously selects the best proxy spacing based on the predetermined proxy count for different points depending on their AABB size. Compared to the other sampling methods in Fig. 3 (a) and (b), the spatial-wise proxy sampling method guarantees an appropriate aspect ratio for proxy spacing and specifies a distinct number of proxies. The pseudo-code of spatial-wise sampling is shown in supplementary materials.
@@
Vertex-based Association. A commonly adopted method for association uses the K-Nearest Neighbor (KNN) [49] method, which connects each data point with its $k$ nearest proxies. This approach provides considerable flexibility in defining the degree of correlation between the proxies and the points, thereby permitting modifications in the number of associations. Despite this, the KNN method requires evaluating the distances between all points and proxies, which adds computational complexity. In KNN using the L-Infinity Norm with $k=8$, every point is linked to 8 proxies positioned at the corners of a cubic proxy within the AABB coordinate sphere. As the proxy is formed through grid sampling, its structural layout can be exploited to create an efficient association method, thereby reducing the need for intricate distance computations. As illustrated in Fig. 4, the initial transformation involves converting the coordinates of the points to their corresponding locations within the proxy grid. Then, the positions of the nearest eight proxies are determined by computing the floor and ceiling for each coordinate. Finally, the compilation of association tuples $as$ is achieved by merging all associations into a list.